OpenStudy (anonymous):

please help & explain how to do it! Solve the compound inequality 60 < 20u < 120. 3 < u < 6 6 > u > 7 6 < u < 3 3 > u < 6

OpenStudy (anonymous):

divide all the way across by 20 and you will have the answer is one step

OpenStudy (anonymous):

answer a) if that was not clear

OpenStudy (anonymous):

\[\frac{60}{20}<\frac{20u}{20}<\frac{120}{20}\] \[3<u<6\]